# Single-Domain Generalized Object Detection by Balancing Domain Diversity and Invariance

### v1

@@ 4.2 Performance Analysis
Results on dusk-rainy conditions. As shown in Tab. 5, our method demonstrates exceptional performance across all categories, particularly in the Bus, Car, and Truck categories. Furthermore, the overall performance improves by 3.1% and 1.7% compared to the CLIP-Gap [60] and Prompt-D [28], the most recent methods, respectively. The robustness of our method in handling challenging environments is also evident. The visual analysis presented in Fig. 4 illustrates the diversity and invariance of the equilibrium domain, which contributes to effective model training.
Methods $\mathcal{L}_{\mathcal{C}}+\mathcal{L}_{\mathcal{H}}$ $\mathcal{L}_{\mathcal{FD}}$ $\mathcal{L}_{\mathcal{WAM}}$ DC DF NS
FR [33] 54.7 32.2 34.0
$\checkmark$ 55.8 37.1 39.3
$\checkmark$ 54.1 37.6 39.4
$\checkmark$ $\checkmark$ 55.9 38.0 39.7
$\checkmark$ 56.5 37.7 40.3
$\checkmark$ $\checkmark$ $\checkmark$ 56.1 39.3 42.0
Table 6: Ablation study. Here, DC, DF, and NS stand for daytime-clear, daytime-foggy and night-sunny, respectively.
4.3 Ablation Study
@@
To assess the influence of each component on the performance of DIDM, we performed an ablation study. The model was trained utilizing the daytime-clear dataset and subsequently evaluated on both the daytime-foggy and night-sunny. Through a comprehensive analysis of the individual contributions of DLM and WAM, we confirmed the efficacy of both modules in the context of object detection. Furthermore, we illustrated the synergistic effect achieved by integrating these two modules, which significantly improves the model’s generalization capabilities and detection accuracy. The conclusive results are detailed in Tab. 6.
@@
The first component is the DLM, as illustrated in Tab. 6. The model’s performance can be significantly improved by utilizing $\mathcal{L}_{\mathcal{H}}$ and $\mathcal{L}_{\mathcal{FD}}$, respectively. This suggests that either constraining the semantic information of domain-specific features or preserving the diversity of these features can enhance the model’s detection capabilities. Furthermore, the model’s performance is further augmented by employing both $\mathcal{L}_{\mathcal{H}}$ and $\mathcal{L}_{\mathcal{FD}}$. This demonstrates that the DLM component effectively enhances the model’s ability to extract semantic information from images. In Tab. 6, we analyze the role of the WAM and observe that the model’s generalization performance improves by 5.5% and 6.3% for daytime-foggy and night-sunny conditions, respectively, compared to the baseline Faster R-CNN [33]. Additionally, the model’s test map achieves a score of 56.5%. This indicates that the WAM enhances the model’s ability to learn domain-invariant features while maintaining a degree of diversity in domain-specific features by incorporating loss weights.
@@
(a) Analysis of $\alpha$ when $\gamma$ is 0.45
(b) Analysis of $\gamma$ when $\alpha$ is 0.45
Figure 5: Hyperparameter Analysis.
4.4 Hyperparameter Analysis
Appropriate hyperparameter settings are essential for optimizing model performance. For the parameters $\alpha$ and $\gamma$ in Eq. 10, we experimented with various hyperparameter configurations to identify the optimal settings. As illustrated in Figs. 5(a) and 5(b), the model demonstrates the best performance when $\alpha$ =0.45 and $\gamma$ = 0.45.
5 Conclusion

### v2

@@ IV-C Performance on DWD datasets
Results on Dusk-Rainy conditions. As shown in Table VI, our method demonstrates exceptional performance across all categories, particularly in the Bus, Car, and Truck categories. Furthermore, the overall performance improves by 2.6% and 3.1% compared to the Prompt-D [15] and UFR [14], the most recent methods, respectively.
Figure 4: Qualitative evaluation results for Real to Artistic datasets are presented. The top row showcases the detection visualizations generated by the Faster R-CNN [1] model, while the bottom row displays the detection visualizations produced by the DIDM.
TABLE VII: DA and DG Results on Real to Artistic Datasets
Methods Clipart Comic watercolor
DA methods
SWDA [36] 38.1 27.4 53.3
I3Net [29] - 30.1 51.5
HTCN [5] 40.3 - -
DBGL [64] 41.6 29.7 53.8
DG methods
FR [1] 25.1 16.8 45.5
NP [65] 35.4 28.9 53.3
Div. [40] 33.7 25.5 52.5
DivAlign [40] 38.9 33.2 57.4
GDD(R101) [66] 40.8 29.7 54.2
FGT(R101) [67] 40.5 30.0 56.6
Ours 34.5 31.7 57.3
Visualization Results The visual analysis presented in Fig. 3 clearly demonstrates that balancing domain diversity and domain invariance is crucial for enhancing the model’s generalization capabilities. The quantitative results indicate that the proposed method significantly outperforms the Faster R-CNN [1] baseline in detection visualization, effectively addressing the issues of misdetection and omission identified in CLIP-Gap [39], with its detection accuracy nearing the level of actual labeling. Notably, the DIDM method consistently exhibits exceptional recognition accuracy in challenging and complex scenarios, such as Night-Rainy conditions, thereby validating its effectiveness in improving the robustness of target detection.
IV-D Real to Artistic task
In this study, we thoroughly evaluate the generalization performance of the DIDM method for art style images. We provide a detailed comparison with the current mainstream domain generalization methods DivAlign [40], NP [65], GDD(R101) [66], and FGT(R101) [67] as well as various domain adaptation methods SWDA [36], I3Net [29], HTCN [5] and DBGL [64]. The results are shown in Table VII. Within the three major domains—Clipart, Comic, and Watercolor, which exhibit significant stylistic differences, the method presented achieves stable performance, effectively validating the model’s strong adaptability to complex distributional biases. Specifically, among the domain generalization (DG) methods, our approach achieves a detection accuracy of 57.3% in the Watercolor domain, which is nearly equivalent to the current optimal DivAlign method (57.4%), while significantly outperforming other DG benchmark methods. Our methods also outperforms some of DA models which require target domain data. Therefore, the effectiveness of the proposed method for enhancing the feature diversity is proven.
Visualization Results We also present the visualization result of the experiment in Fig. 4, where the first and second rows show the detection results without transfer learning and our methods, respectively. Obviously, our method achieves better detection performance.
TABLE VIII: Single-Domain Generalization Object Detection Results. We trained the Model Using One dataset as the Source, while Also Assessing Its Generalization Across the Other Three Domains.
Source Target Methods Ref Cityscaps Foggy Cityscapes Rainy Cityscapes BDD-100K
Foggy Cityscapes [61] FR [1] NeurIPS’15 33.3 - 43.7 13.9
FACT [68] CVPR’21 30.0 - 38.7 20.2
FSDR [69] CVPR’22 31.3 - 40.8 20.4
MAD [70] CVPR’23 41.3 - 43.3 24.4
Ours - 44.5 - 50.2 24.1
BDD-100K [55] FR [1] NeurIPS’15 37.4 26.9 45.2 -
FACT [68] CVPR’21 32.4 24.3 33.9 -
FSDR [69] CVPR’22 32.4 27.8 34.7 -
MAD [70] CVPR’23 36.4 30.3 36.1 -
Ours - 38.2 31.4 45.4 -
TABLE IX: The Quantitative Results (%) from Cityscapes to BDD-100K.
Methods Bus Bike Car Motor Person Rider Truck mAP
FR [1] 33.5 26.8 41.3 13.0 16.2 11.2 23.6 23.7
SW [48] 22.2 21.3 36.7 20.9 25.6 23.1 18.8 24.1
ISW [43] 21.6 20.9 35.2 17.8 22.7 22.4 16.6 22.5
IBN-Net [42] 19.2 14.9 31.9 13.7 21.4 19.3 13.4 19.1
IterNorm [44] 21.6 21.2 36.2 20.9 23.4 10.6 18.0 23.7
CDSD [45] 20.5 22.9 33.8 14.7 18.5 23.6 18.2 21.7
SHADE [49] 19.0 25.1 36.8 18.4 24.1 24.9 19.8 24.0
SRCD [50] 21.5 24.8 38.7 19.0 25.7 28.4 23.1 25.9
MAD [70] - - - - - - - 28.0
Ours 40.4 33.2 56.0 19.2 20.3 19.4 30.4 31.3
IV-E Diversified Weather datasets
Here, we compare DIDM with recent methods object detection such as FACT [68], FSDR [69], and MAD [70], as well as with four feature-based normalization techniques: SW [48], IBN-Net [42], IterNorm [44], and ISW [43]. This comparison aims to further validate the applicability and effectiveness of DIDM across various domains.
Results on Diversified Weather. In this study, we train the model using the only one datasets, and evaluate its performance on the remaining datasets to thoroughly assess the generalization capability of the DIDM method. As shown in Table VIII, our approach exhibits outstanding performance across datasets with varying weather conditions. This highlights the effectiveness of balancing the diversity of domain-invariant features with domain-specific features during the learning process.
Results on BDD-100K. Table IX presents a performance comparison of the method proposed in this paper with several contemporary DGOD methods across various categories. The results indicate that the DIDM method demonstrates exceptional generalization performance in all categories, particularly in the bus, car, and truck categories, where it significantly surpasses the other methods. This is a strong indication that the approach outlined in this paper achieves superior generalization and accuracy in the target detection task.
TABLE X: Car Class Detection Results (%) from Sim10k Dataset to Cityscapes and BDD100K Datasets.
Methods Ref Cityscapes BDD-100K
FR [1] NeurIPS’15 34.0 30.0
IBN-Net [42] ECCV’18 33.2 25.7
SW [48] ICCV’19 34.5 30.0
IterNorm [44] CVPR’19 34.3 25.7
ISW [43] CVPR’21 40.4 28.5
CDSD [45] CVPR’22 35.2 27.4
SHADE [49] CVPR’22 40.9 30.3
SRCD [40] TNNLS’24 43.0 31.6
Ours - 51.6 39.8
IV-F Artificial to Real datasets
Here, we compare DIDM with recent ISW [43], CDSD [45], SRCD [40] and four feature-based normalization methods. As shown in Table X, the results of car category detection on artificial-to-real datasets demonstrate that the proposed method exhibits exceptional performance on both the Cityscapes and BDD-100K datasets. It significantly outperforms the previous best SRCD [40] method, which achieved detection results of 43.0% and 31.6%, with our method achieving 51.6% and 39.8%, respectively. This fully validates its efficient migration capability and robustness in synthetic-to-real scenarios within the cross-domain detection task.
TABLE XI: The Results of Our Proposed Diversity Invariant Detection Model (DIDM) Ablation Experiment are Analyzed. The Losses Introduced in this Paper are Progressively Incorporated into the Baseline Network Faster R-CNN [1], and the Contribution of Each Module to Target Detection is Examined.
Methods $\mathcal{L}_{\mathcal{C}}+\mathcal{L}_{\mathcal{H}}$ $\mathcal{L}_{\mathcal{FD}}$ $\mathcal{L}_{\mathcal{WAM}}$ Daytime - Clear Night - Sunny Dusk - Rainy Night - Rainy Daytime - Foggy
FR [1] 57.3 38.1 32.3 14.6 34.5
$\checkmark$ 56.9 41.9 34.4 17.1 38.6
$\checkmark$ 58.0 41.8 33.3 16.8 38.4
$\checkmark$ $\checkmark$ 57.7 42.4 35.7 19.2 39.4
$\checkmark$ 57.9 42.2 35.5 18.7 39.4
$\checkmark$ $\checkmark$ $\checkmark$ 57.7 43.4 36.3 20.4 40.1
V Model Analysis and Discussion
In this section, we analyze and discuss our model with several different experiment to show its characteristics and effectiveness. An in-depth insight into our model is presented.
V-A Ablation Study
@@
To assess the influence of each component on the performance of DIDM, we performed an ablation study. The model was trained using the daytime-clear datasets and subsequently evaluated on the remaining four datasets: night-sunny, dusk-rainy, night-rainy, and daytime-foggy. Through a comprehensive analysis of the individual contributions of DLM and WAM, we confirmed the efficacy of both modules in the context of object detection. Furthermore, we illustrated the synergistic effect achieved by integrating these two modules, which significantly improves the model’s generalization capabilities and detection accuracy. The conclusive results are detailed in Table XI.
Figure 5: Hyperparameter Analysis. (a) Analysis of $\tau$. (b) Analysis of $\beta$. (c) Analysis of the Choice of Hyperparameter $\alpha$.
@@
The first component is the DLM, as illustrated in Table XI. The model’s performance can be significantly improved by utilizing $\mathcal{L}_{\mathcal{H}}$ and $\mathcal{L}_{\mathcal{FD}}$, respectively. This suggests that either constraining the semantic information of domain-specific features or preserving the diversity of these features can enhance the model’s detection capabilities. Furthermore, the model’s performance is further augmented by employing both $\mathcal{L}_{\mathcal{H}}$ and $\mathcal{L}_{\mathcal{FD}}$. This demonstrates that the DLM component effectively enhances the model’s ability to extract semantic information from images. In Table XI, we analyze the role of the WAM and observe that the model’s generalization performance improves by 4.9% and 4.3% for daytime-foggy and night-sunny conditions, respectively, compared to the baseline Faster R-CNN [1]. Additionally, the model’s test MAP achieves a score of 57.9%. This indicates that the WAM enhances the model’s ability to learn domain-invariant features while maintaining a degree of diversity in domain-specific features by incorporating loss weights.
@@
Figure 6: Analysis of Differences in the Distribution of Features. Where Foggy represents the source domain, Cityspace and Rainy represent different target domains, respectively
Figure 7: A-distance Result Display. Where A represents the absence of FD and $\beta$; B indicates the absence of FD; C signifies the absence of $\beta$; and D denotes full DIDM.
V-B Hyperparameter Analysis
Appropriate hyperparameter settings are essential for optimizing model performance.
First, in order to effectively differentiate between individual domain-specific features while retaining meaningful domain-specific information, the selection of hyperparameter $\tau$ in Eq. 8 is crucial. As shown in Fig. 5a, the model demonstrates the best generalization performance across multiple datasets when $\tau$ = 0.1. This clearly demonstrates that the appropriate hyperparameter A can effectively balance feature differentiation and information retention, thereby enhancing the model’s adaptability and robustness across various domains.
Second, to verify the effectiveness of the dynamic weighting parameter $\beta$ in Eq. 11, we conducted a comparative experiment. One group utilized parameter $\beta$ when the features were aligned, while the other group did not. The experimental results, illustrated in Fig. 5b, indicate that the use of weighting parameter $\beta$ effectively prevents the over-alignment of features, thereby better maintaining feature diversity. This characteristic significantly enhances the model’s generalization ability and robustness across different datasets, clearly demonstrating the crucial role of parameter $\beta$ in optimizing the feature alignment process.
Last, as shown in Fig. 5c, we experimented with the hyperparameters $\alpha$ in Eq. 13. The results indicate that the selection of hyperparameters significantly impacts model performance, with the model achieving optimal results across different datasets when parameter $\alpha$ = 0.45 is utilized.
V-C Analysis of Differences in the Distribution of Features
The visualization results of the feature distributions for the source and target domains, as illustrated in Fig. 6 and Fig. 7, demonstrate significant overlapping regions between the feature distributions of the two domains. This finding confirms that the DIDM can effectively balance domain-invariant and domain-specific features, thereby reducing inter-domain distribution differences while preserving the discriminative features in cross-domain tasks. The A-distance for Cityspace from Foggy and Rainy conditions is 0.58 and 0.73, respectively.
VI Conclusion